# From Strangers to Assistants: Fast Desire Alignment for Embodied Agent-User Adaptation

### v1

@@ 5.2 Baselines & Ablations
We compare FAMER against three baselines and evaluate its components through three ablated variants.
CoELA zhang2023building : An LLM-based multiagent cooperation framework that includes perception, communication, planning, memory, and execution modules. In its original form, CoELA assumes full observability of goals. To adapt it to our setting, we modify its prompting so that the agent is only aware of the potential goal set and the number of target goals.
ProAgent zhang2024proagent : A proactive LLM-based agent framework designed for collaboration in fixed-goal tasks. It lacks mechanisms for handling goal uncertainty or communication. We extend ProAgent by adding cross-episode memory and explicit feedback to support adaptation to latent user desires.
MHP: An MCTS-based Hierarchical goal-sampling Planner adapted from the Watch-and-Help Challenge puig2020watch . We introduce subgoal sampling to handle uncertain goals and maintain a cross-episode success memory. Like ProAgent, MHP does not support communication.
@@
Figure 4: Quantitative results in Snack-M and Snack-L.
Figure 5: Quantitative results in Table-M and Table-L.
5.3 Quantitative Results
@@
We evaluate performance using the three metrics on both Snack and Table tasks at two difficulty levels: Medium and Large. Results for the Snack-M and Snack-L tasks are shown in Figure 4, while those for Table-M and Table-L are presented in Figure 5. For each method, the three adjacent bars represent performance across three consecutive episodes with the same user.
@@
From the figures, it is evident that FAMER consistently outperforms all three baselines across all metrics. In terms of score, FAMER achieves the maximum value of 1.0 in nearly every episode, indicating that it successfully completes all desired goals while avoiding incorrect ones. CoELA performs second best but falls short due to its reliance on long-context LLM prompting alone, which leads to occasional misinterpretation of user desires. This limitation will be further illustrated in the qualitative analysis section. MHP and ProAgent perform the worst, as they lack communication capabilities and rely solely on trial-and-error to identify goals. Such an inefficient process often incurs penalties. Notably, their performance gradually improves across episodes, reflecting slow adaptation to latent user desires through repeated interactions.
FAMER also demonstrates superior efficiency in execution, requiring significantly fewer environment steps to complete tasks. MHP and ProAgent often exhaust the full episode length without completing the goals. This improvement is a direct result of FAMER’s integrated components: desire-centered reasoning rapidly narrows the goal space, efficient communication accelerates goal confirmation, and key information extraction reduces unnecessary exploration by reusing known object-location pairs.
@@
In terms of communication cost, FAMER significantly outperforms CoELA, as shown in Figures 4 and 5. This efficiency stems from FAMER’s reflection-based communication strategy, which avoids repeated or redundant questions. In contrast, CoELA frequently issues similar or vague queries due to its lack of explicit goal-tracking mechanisms.
@@
Ablation Study We further evaluate the contribution of each FAMER component through ablation studies on the Snack-M task. As shown in Figure 6, all three ablated variants show performance degradation across the evaluated metrics. Among them, FAMER w/o KeyInfo maintains the same score as FAMER but requires more steps, demonstrating that the KeyInfo Extraction module improves efficiency by reducing unnecessary exploration, rather than directly influencing goal correctness.
@@
A significant performance drop is observed in both FAMER w/o Desire and FAMER w/o EC, highlighting that desire modeling and efficient communication are central to agent-user adaptation. Without goal confirmation, desire inference, and goal-aligned action filtering, the agent struggles to accurately interpret user intent and often executes incorrect or inefficient actions.
Figure 6: Ablation results in Snack-M.
5.4 Qualitative Analysis
@@ 6 Conclusion
We address the critical problem of adapting embodied agents to unfamiliar users with implicit values and desires, which is a key challenge for real-world deployment of assistive AI. To facilitate development and evaluation in this setting, we introduce HA-Desire, a novel 3D simulation environment featuring value-driven proxy users, natural language communication, and object-rich household tasks. Unlike prior benchmarks, HA-Desire captures the complexity of real-world assistance by simulating ambiguous goal specifications and indirect, human-like communication.
@@
Extensive experiments on two representative tasks (Snack and Table) at varying difficulty levels show that FAMER consistently outperforms strong baselines in task success rate, execution efficiency, and communication cost. Ablation studies confirm the significance of each component, particularly the central role of desire modeling in achieving user-aligned behavior. Qualitative analyses further highlight FAMER’s advantages in intent inference, planning precision, and communication efficiency.
@@ Appendix A Environment Details
HA-Desire is built upon VirtualHome [28], a widely adopted testbed for embodied multi-agent cooperation. In this work, we modify the simulation to specifically address the challenge of embodied agent-user adaptation with a focus on desire alignment. In the following sections, we provide further details on the human user agent and the tasks designed to evaluate this problem.
@@ A.2 Tasks
As described in Section 5, we evaluate our method, along with baselines and ablations, on two tasks: Prepare Afternoon Snack and Set Up Dinner Table. The detailed settings for these tasks are provided below.
Snack. This task involves ten potential goals: cupcake, wine, milk, cereal, chips, apple, juice, pudding, creamybuns, and chocolatesyrup. The value space consists of five dimensions: Hungry, Thirsty, SweetTooth, Fruitarian, and Alcoholic, each of which takes one of three discrete levels: Not, Somewhat, or Very. The human user randomly samples values for these dimensions and then uses a language model to generate the corresponding goal set. The Snack-M level represents the medium difficulty, with 2 goals and a maximum of 200 steps. The Snack-L level represents the large difficulty, with 4 goals and a maximum of 300 steps. Performance is evaluated using three metrics: score, step count, and communication cost, as described in Section 5.
@@ B.1 Computing Resource
The experiments were conducted on a workstation equipped with an NVIDIA GeForce RTX 4090 GPU and an Intel Core i9-13900K CPU. The large vision-language model used in this study is the GPT-4o.
B.2 Baselines
@@
B.3 Qualitative Analysis
Figure 7 provides a qualitative comparison between CoELA and FAMER. We present more details about the behavior of both methods in the Snack-L task, where the agent must infer 4 goals from a set of ten potential options: cupcake, wine, milk, cereal, chips, apple, juice, pudding, creamybuns, and chocolatesyrup.
For CoELA, the ground truth goal set is: creamybuns, milk, juice, chips. As shown in Figure 7, the CoELA agent makes several mistakes during the task. Initially, the agent incorrectly interprets the user’s message and grabs apple (step 1). It then retrieves chips (step 2) and places both apple and chips on the coffee table (step 3). However, the agent continues to confuse apple as a valid goal, repeatedly placing it on the table in step 4. After some dialogue with the user, it begins searching for the other correct goal objects, successfully finding creamybuns (step 5), but erroneously selects cupcake (step 6) as a goal. Eventually, the agent locates milk and places all the goal objects on the table, but with two incorrect items (apple and cupcake) (steps 7 & 8).
For FAMER, the ground truth goal set is: cereal, creamybuns, cupcake, apple. As shown in the figure, the FAMER agent efficiently infers the user’s desires through a series of concise interactions. First, it grabs creamybuns (step 1) and cereal (step 2) and places them on the coffee table (step 3), since it can hold only two objects at a time. It then searches for apple (step 4) and cupcake (step 5), placing both objects on the table (step 6) and successfully completing the task.

### v2

@@ 5.2 Baselines & Ablations
We compare FAMER against three baselines and three ablated variants.
CoELA Zhang et al. (2024b): An LLM-based multiagent cooperation framework that includes perception, communication, planning, memory, and execution modules. In its original form, CoELA assumes full observability of goals. To adapt it to our setting, we modify its prompting so that the agent is only aware of the potential goal set and the number of target goals.
ProAgent Zhang et al. (2024a): A proactive LLM-based agent framework designed for collaboration in fixed-goal tasks. It lacks mechanisms for handling goal uncertainty or communication. We extend ProAgent by adding cross-episode memory to support adaptation to latent user desires.
MHP: An MCTS-based Hierarchical goal-sampling Planner adapted from the Watch-and-Help Challenge Puig et al. (2021). We introduce subgoal sampling to handle uncertain goals and maintain a cross-episode success memory. Like ProAgent, MHP does not support communication.
@@
Figure 4: Quantitative results in Snack-M and Snack-L. Ep. denotes Episode, M denotes Medium, and L denotes Large, which applies to all subsequent figures and tables. FAMER achieves a perfect score by the third episode and outperforms baselines in both task score and communication efficiency.
Figure 5: Quantitative results in Table-M and Table-L. FAMER achieves higher scores and superior communication efficiency compared to all baselines.
5.3 Experimental Results
@@
We evaluate performance using the two metrics on both Snack and Table tasks at two difficulty levels: Medium and Large. Results for the Snack-M and Snack-L tasks are shown in Figure 4, while those for Table-M and Table-L are presented in Figure 5. For each method, the three adjacent bars represent performance across three consecutive episodes with the same user.
@@
From the figures, it is evident that FAMER consistently outperforms all three baselines across all metrics. In terms of score, FAMER achieves the maximum value of 1.0 in the third episode, indicating that it successfully infers the desired goals of the human user within only 3 episodes and completes all of them. CoELA performs second best but falls short due to its reliance on long-context LLM prompting alone, which leads to occasional misinterpretation of user desires. This limitation will be further illustrated in the qualitative analysis in Appendix B.4. MHP and ProAgent perform the worst, as they lack communication capabilities and rely solely on trial-and-error to identify goals. Such an inefficient process often incurs penalties. Notably, their performance gradually improves across episodes, reflecting slow adaptation to latent user desires through repeated interactions.
@@
In terms of communication efficiency, FAMER significantly outperforms CoELA, as shown in Figures 4 and 5. This efficiency stems from FAMER’s reflection-based communication strategy, which avoids repeated or redundant questions. In contrast, CoELA frequently issues similar or vague queries due to its lack of explicit goal-tracking mechanisms.
Figure 6: Ablation results on Snack-M. Removing any component of FAMER degrades performance in both scores and communication efficiency. w/o KeyInfo is least affected, still reaching full score by the third episode. In contrast, w/o EC and w/o Desire both cause notable score drops, with w/o EC also sharply raising communication cost. These results confirm the importance of all three modules.
5.4 Ablation Study
@@
We further evaluate the contribution of each FAMER component through ablation studies on the Snack-M task. As shown in Figure 6, all three ablated variants exhibit performance degradation across the evaluated metrics. Among them, FAMER w/o KeyInfo is able to eventually achieve a full score in the third episode, similar to the full model, but its scores in the first two episodes are lower and it incurs slightly higher communication costs. This indicates that the Key Information Extraction module primarily improves efficiency by reducing unnecessary exploration.
Figure 7: Impact of maximum episode length on average score in Snack-M.
@@
In contrast, substantial performance drops are observed in FAMER w/o Desire and w/o EC, underscoring the central roles of desire modeling and efficient communication in agent-human adaptation. Without goal confirmation, desire inference, and goal-aligned action filtering, the agent struggles to correctly interpret user intent, leading to incorrect or inefficient actions. Moreover, the sharp increase in communication cost for FAMER w/o EC highlights that reflection-based communication is crucial for minimizing redundant messages and maintaining efficiency.
Impact of Episode Length on Score. We evaluate how maximum episode length affects performance on Snack-M by varying the limit from 20 to 100 steps and averaging scores across three episodes (Figure 7). When the limit is below 30, all methods score zero, as no goal can be completed. With more steps, FAMER and CoELA improve in-episode due to their ability to communicate, while FAMER reaches a perfect score once the limit exceeds 90 steps, showing it can fully adapt in the first episode given enough steps. In contrast, ProAgent and MHP may decline as episode length grows, since they rely on trial-and-error for adaptation, which leads to more wrong-goal penalties.
Human Reveal Goals. We also test a setting where the human user directly reveals the goals in each episode, removing the need for goal inference. Results on Snack-M (Table 2) show all methods improve, but FAMER still leads. Its Key Information Extraction module stores goal-related details like object positions across episodes, avoiding redundant exploration, while baselines such as CoELA must still search extensively in each episode. Combined with goal-oriented planning that prunes irrelevant actions, this allows FAMER to outperform baselines even when goals are explicitly given.
5.5 Human Study
To evaluate how FAMER performs with real human users, we recruited 8 participants to serve as users in two scenarios: Snack-M and Table-M. Each participant was randomly assigned a set of value attributes and asked to act as the human user in HA-Desire, communicating with the agents and evaluating whether they successfully satisfied the assigned values.
For each task, participants interacted with four agents: CoELA, FAMER w/o Desire, FAMER w/o EC, and FAMER, in random order. Each agent was tested for 3 episodes, following the same protocol as in the simulation setting. After each session, participants rated the assigned agent on a 7-point Likert scale with respect to three dimensions: (1) Satisfaction: I am satisfied with the overall performance of the agent. (2) Helpfulness: The agent helped me obtain what I wanted. (3) Communication Efficiency: The agent communicated efficiently without asking redundant or irrelevant questions. Each participant evaluated all four agents across both tasks for three episodes, resulting in a total of 192 episodes of human-agent interactions.
Results are presented in Table 2. FAMER consistently achieved the highest ratings across all three criteria, further supporting its effectiveness in real human-agent adaptation scenarios.
Table 1: Score results across three episodes in the Human Reveal Goals setting on Snack-M.
Method Ep. 1 Ep. 2 Ep. 3
CoELA 0.67 0.83 0.83
MHP 0.42 0.33 0.42
ProAgent 0.50 0.50 0.58
FAMER 0.67 1.00 1.00
Table 2: Human study results on Snack-M and Table-L. Participants rated Satisfaction, Helpfulness, and Communication Efficiency on a 7-point Likert scale.
Method Satisfaction Helpfulness Comm Efficiency
[…]
@@ 6 Conclusion
We address the critical problem of adapting embodied agents to unfamiliar human users with implicit values and desires, which is a key challenge for real-world deployment of assistive AI. To facilitate development and evaluation in this setting, we introduce HA-Desire, a novel 3D simulation environment featuring value-driven proxy users, natural language communication, and object-rich household tasks. Unlike prior benchmarks, HA-Desire captures the complexity of real-world assistance by simulating ambiguous goal specifications and indirect, human-like communication.
@@
Extensive experiments on two representative tasks at varying difficulty levels show that FAMER consistently outperforms strong baselines in task execution and communication cost. Ablation studies confirm the significance of each component, particularly the central role of desire modeling in achieving user-aligned behavior. Additional analysis, including the human-revealed-goal setting, the impact of episode length, and a human-subject study, further validates the robustness of our approach.
Future work includes deploying FAMER on real robotic platforms to assess its effectiveness in real-world physical environments, and exploring post-training strategies to further align large language models with human values and preferences.
@@ Appendix A Environment Details
HA-Desire is built upon VirtualHome Puig et al. (2018), a widely adopted testbed for embodied multi-agent cooperation. In this work, we modify the simulation to specifically address the challenge of embodied agent-human adaptation with a focus on desire alignment. In the following sections, we provide further details on the human user agent and the tasks designed to evaluate this problem.
@@ A.2 Tasks
As described in Section 5, we evaluate our method, along with baselines and ablations, on two tasks: Prepare Afternoon Snack and Set Up Dinner Table. The details of these tasks are provided below.
Snack. This task involves ten potential goals: cupcake, wine, milk, cereal, chips, apple, juice, pudding, creamybuns, and chocolatesyrup. The value space consists of five dimensions: Hungry, Thirsty, SweetTooth, Fruitarian, and Alcoholic, each of which takes one of three discrete levels: Not, Somewhat, or Very. The human user randomly samples values for these dimensions and then uses a language model to generate the corresponding goal set. The Snack-M level represents the medium difficulty, with 2 goals and a maximum of 60 steps. The Snack-L level represents the large difficulty, with 4 goals and a maximum of 120 steps. Performance is evaluated using two metrics: score and communication cost, as described in Section 5.
@@ B.1 Computing Resource
The experiments were conducted on a workstation equipped with an NVIDIA GeForce RTX 4090 GPU and an Intel Core i9-13900K CPU. The large vision-language model used in this study is GPT-4o.
B.2 FAMER Implementation
The perception module of FAMER in HA-Desire follows the design of CoELA Zhang et al. (2024b). It employs a Mask-RCNN to generate segmentation masks from RGB images, then combines them with depth information to build 3D point clouds of objects. From these, the agent extracts high-level information such as the position of key objects and constructs a structured semantic map for downstream reasoning and planning.
The memory module of FAMER maintains several types of information, as illustrated in Figure 3: Confirmed Goals, KeyInfo Context, Task Progress, Previously Achieved Goals, and Action & Dialogue History. The first four categories represent compact summaries of task state and goal inference, and thus grow slowly during interaction; each typically contains fewer than 20 entries, making it feasible to store them entirely. In contrast, Action & Dialogue History grows linearly with the number of steps. To manage this, we retain only the latest 10 entries in the memory context. Since the key information context preserves important earlier details, essential knowledge from prior interactions is not lost.
B.3 Baselines
@@
B.4 Qualitative Analysis